### v3

@@ Abstract
Recently, Large Language Models (LLMs) have dominated much of the artificial intelligence scene with their ability to process and generate natural languages. However, the majority of LLM research and development remains English-centric, leaving low-resource languages such as those in the SouthEast Asian (SEA) region under-represented. To address this representation gap, we introduce Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, two cutting-edge multilingual LLMs designed for SEA languages. The SEA-LION family of LLMs supports 11 SEA languages, namely English, Chinese, Indonesian, Vietnamese, Malay, Thai, Burmese, Lao, Filipino, Tamil, and Khmer. Our work leverages large-scale multilingual continued pre-training with a comprehensive post-training regime involving multiple stages of instruction fine-tuning, alignment, and model merging. Evaluation results on multilingual benchmarks show that our models achieve state-of-the-art performance across LLMs supporting SEA languages. We open-source the models111[LINK_HIDDEN] to benefit the wider SEA community.
@@ 1 Introduction
Large language models (LLMs) have significantly transformed the field of natural language processing, achieving remarkable performance in text generation, summarization and sentiment analysis (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Rivière et al., 2024; Zhang et al., 2024b; Yeo et al., 2024).
@@
Despite their impressive capabilities, most LLMs remain heavily English-centric (Wendler et al., 2024; Zhong et al., 2024). Unfortunately, this situation has led LLMs in regions with many under-represented languages such as Southeast Asia (SEA) to suffer. Languages with lower resources, such as Filipino, Lao, Burmese and Khmer in the SEA region, are not supported by many open-source English-centric LLMs. This underscores the need to bridge the resource and representation gap between English and SEA languages.
@@
Recently, there have been many attempts to create multilingual LLMs in an open-source manner, e.g., BLOOM (Scao et al., 2022), a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks. Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
@@
In this paper, we address the issues by proposing a robust open-source Southeast Asian model with data transparency for reproducibility, namely SEA-LION – a family of LLMs continued pre-trained (CPT) and fine-tuned on Llama-3.1-8B-Instruct for Llama-SEA-LION-8B-IT and Gemma-2-9B for Gemma-SEA-LION-9B-IT with a focus on SEA languages. To tackle the performance problem, we utilize 200 billion English, code and SEA languages tokens as well as 16.8 million English and SEA languages instruction and answer pairs for CPT and post-training steps respectively, to achieve a significant improvement in SEA languages. In order to allow our models to be used by everyone without restrictions, we release our models under a fully open MIT  [^3] https://mit-license.org/. We benchmark our models against the SEA-HELM(Susanto et al., 2025) and Open LLM Leaderboard [^4] Open LLM Leaderboard with other LLMs of similar sizes in Southeast Asia like Sailor 2 (Team, 2024) and SeaLLMs 3 (Zhang et al., 2024a), where our models achieve state-of-the-art performances. We summarize the contribution of our paper as follows.
@@
    We present a reproducible dataset development process, covering sourcing and the model training process. Upon the acceptance of this paper, we will release our training artifacts, including the training dataset, training scripts, training checkpoints, and fine-tuned models, including Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, to provide strong baselines, promote reproducibility, and enable future research on applications that require SEA-specific knowledge.
@@ 2.1 Pre-training data
A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021). To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English, and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to Table 6.
@@ 2.2 CPT process
Training setup. Following previous works Dou et al. (2024), we use BPE-Dropout (Provilkov et al., 2020) to increase the performance and robustness of the training. We use a Warmup-Stable-Decay (WSD) (Hu et al., 2024) scheduler with warm-up and cooldown phases each representing 10% of the entire training budget. We use the AdamW (Loshchilov and Hutter, 2019) optimiser with the maximum learning rate (LR) set to $1e^{-5}$} and the final LR after cooldown is $1e^{-7}$}. Following Wortsman et al. (2024), we set epsilon to $1e^{-15}$}. We use Composer (Team, 2021) and LLM Foundry (Team, 2022) for distributed training using Fully Sharded Data Parallel (Zhao et al., 2023) on a cluster of eight nodes of the p5.48xlarge instance from Amazon Web Services (AWS). The total training duration was approximately 6 days and 10 days for the Llama 3.1 and Gemma 2 models, respectively. In this paper, we refer to the post-CPT models as Llama-SEA-LION-8B and Gemma-SEA-LION-9B for the Llama 3.1 and Gemma 2 continued pre-trained models respectively.
@@ 3.2 Post-training process
We use LLaMaFactory (Zheng et al., 2024b) with DeepSpeed (Rasley et al., 2020) for all Instruction Fine Tuning (IFT) and alignment steps. All IFT stages are performed using full model fine-tuning where the models are from the previous step (Section 2.2) and existing models. We use MergeKit (Goddard et al., 2024) with a value of 1 for weight and density parameters for all merge steps. Models selected for merging are selected empirically, based on the openness of model licenses, the suitability for merging and performance.
@@ 3.2.1 Llama-SEA-LION-8B-IT
Second merge In order to mitigate catastrophic forgetting due to the fine-tuning process (Alexandrov et al., 2024), we performed the second round of merge by merging top-performing instruction-tuned models that share the Llama 3.1 lineage. We merge the original Llama-3.1-8B-Instruct, Llama3-8B-SEA-LION-v2.1-Instruct (SEA-LION Team, 2024), and SuperNova-Lite (Arcee-AI, 2024) into Merge-1-Llama using the Consensus TA (Wang et al., 2024b; Ilharco et al., 2023) merge method. We refer to the model at this stage as Merge-2-Llama.
@@
Final merge Lastly, we perform a merge using the DELLA-Linear merge. With the original Llama-3.1-8B-Instruct model as the base for merging, we merge in Merge-2-Llama and Aligned-SimPO-Llama to produce the final model, Llama-SEA-LION-9B-IT.
@@ 4.2 Evaluation Benchmarks
SEA Benchmarks. We evaluated the multilingual performance of each LLM using the SEA-HELM Leaderboard (Leong et al., 2023; Susanto et al., 2025). We selected SEA-HELM because the design choice of this benchmark reflects the performance of SEA culture and knowledge the most compared with other existing benchmarks DAMO-NLP-SG (2024); Lovenia et al. (2024); Wang et al. (2024a). We also evaluate on a wide-range SEA coverage language benchmark called SEACrowd Lovenia et al. (2024). This benchmark consists of all SEA languages for natural language understanding and generation datasets. However, due to maintenance reasons, we cannot reproduce the NLG benchmark of SEACrowd. Therefore, we experiment only with the NLU benchmark (zero-shot), which has 131 data subsets, 7 tasks, and 31 SEA indigenous languages.

### v4

@@ Abstract
Recently, Large Language Models (LLMs) have dominated much of the artificial intelligence scene with their ability to process and generate natural languages. However, the majority of LLM research and development remains English-centric, leaving low-resource languages such as those in the Southeast Asian (SEA) region underrepresented. To address this representation gap, we introduce Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, two cutting-edge multilingual LLMs designed for SEA languages. The SEA-LION family of LLMs supports 11 SEA languages, namely English, Chinese, Indonesian, Vietnamese, Malay, Thai, Burmese, Lao, Filipino, Tamil, and Khmer. Our work leverages large-scale multilingual continued pre-training with a comprehensive post-training regime involving multiple stages of instruction fine-tuning, alignment, and model merging. Evaluation results on multilingual benchmarks show that our models achieve state-of-the-art performance across LLMs supporting SEA languages. We open-source the models  [^1] SEA-LION Models Collection to benefit the wider SEA community.
@@ 1 Introduction
Large language models (LLMs) have significantly transformed the field of natural language processing, achieving remarkable performance in text generation, summarization and sentiment analysis (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Rivière et al., 2024; Zhang et al., 2024b; Yeo et al., 2024). Despite their impressive capabilities, most LLMs remain heavily English-centric (Wendler et al., 2024; Zhong et al., 2024). Unfortunately, this situation has led LLMs in regions with many under-represented languages such as Southeast Asia (SEA) to suffer. Languages with lower resources, such as Filipino, Lao, Burmese and Khmer in the SEA region, are not supported by many open-source English-centric LLMs. This underscores the need to bridge the resource and representation gap between English and SEA languages.
@@
Recently, there have been many attempts to create multilingual LLMs in an open-source manner, e.g., BLOOM (Scao et al., 2022), a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., when considering evaluation benchmarks formulated from English datasets. However, we observe that the performance degrades on SEA-specific benchmarks. Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
@@
In this paper, we address the issues by proposing a robust open-source Southeast Asian model with data transparency for reproducibility, namely SEA-LION – a family of LLMs continued pre-trained (CPT) and fine-tuned on Llama-3.1-8B-Instruct for Llama-SEA-LION-8B-IT and Gemma-2-9B for Gemma-SEA-LION-9B-IT with a focus on SEA languages. To tackle the performance problem, we utilize 200 billion English, code, and SEA languages tokens as well as 16.8 million English and SEA languages instruction and answer pairs for CPT and post-training steps, respectively, to achieve a significant improvement in SEA languages. In order to allow our models to be used by everyone without restrictions, we release our models under the fully open MIT license. We benchmark our models against the SEA-HELM(Susanto et al., 2025) and Open LLM Leaderboard [^3] Open LLM Leaderboard with other LLMs of similar sizes in Southeast Asia like Sailor 2 (Team, 2024) and SeaLLMs 3 (Zhang et al., 2024a), where our models achieve state-of-the-art performances.
We summarize the contribution of our paper as follows.
@@
    We present a reproducible dataset development process, covering sourcing and the model training process. We release our training artifacts, including the training dataset, training scripts, training checkpoints, and fine-tuned models, including Llama-SEA-LION-8B-IT and Gemma-SEA-LION-9B-IT, to provide strong baselines, promote reproducibility, and enable future research on applications that require SEA-specific knowledge  [^4] Please visit https://huggingface.co/aisingapore for all artifacts in this paper, including training data and other versions of SEA-LION.
@@ 2.1 Pre-training data
A document is retained if the language code reported in its metadata matches that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021). To determine the optimal dataset ratio between SEA languages, code, and English for the CPT process, we conduct a series of small-scale CPT experiments, each with a training budget of 10 billion tokens and varying proportions of English, code, and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English, and 20% code tokens for a budget of 200 billion tokens. For a detailed breakdown of the token count by languages, please refer to Table 6.
@@ 2.2 CPT process
Training setup. Following previous works Dou et al. (2024), we use BPE-Dropout (Provilkov et al., 2020) to increase the performance and robustness of the training. We use a Warmup-Stable-Decay (WSD) (Hu et al., 2024) scheduler with warm-up and cooldown phases each representing 10% of the entire training budget. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with the maximum learning rate (LR) set to $1e^{-5}$} and the final LR after cooldown is $1e^{-7}$}. Following Wortsman et al. (2024), we set epsilon to $1e^{-15}$}. We use Composer (Team, 2021) and LLM Foundry (Team, 2022) for distributed training using Fully Sharded Data Parallel (Zhao et al., 2023) on a cluster of eight nodes of the p5.48xlarge instance from Amazon Web Services (AWS). The total training duration was approximately 6 days and 10 days for the Llama 3.1 and Gemma 2 models, respectively. In this paper, we refer to the post-CPT models as Llama-SEA-LION-8B and Gemma-SEA-LION-9B for the Llama 3.1 and Gemma 2 continued pre-trained models, respectively.
@@ 3.2 Post-training process
We use LLaMaFactory (Zheng et al., 2024b) with DeepSpeed (Rasley et al., 2020) for all Instruction Fine Tuning (IFT) and alignment steps. All IFT stages are performed using full model fine-tuning, where the models are from the previous step (Section 2.2) and existing models. We use MergeKit (Goddard et al., 2024) with a value of 1 for weight and density parameters for all merge steps. Models selected for merging are selected empirically, based on the openness of model licenses, the suitability for merging and performance.
@@ 3.2.1 Llama-SEA-LION-8B-IT
Second merge In order to mitigate catastrophic forgetting due to the fine-tuning process (Alexandrov et al., 2024), we performed the second round of merging by merging top-performing instruction-tuned models that share the Llama 3.1 lineage. We merge the original Llama-3.1-8B-Instruct, Llama3-8B-SEA-LION-v2.1-Instruct (SEA-LION Team, 2024), and SuperNova-Lite (Arcee-AI, 2024) into Merge-1-Llama using the Consensus TA (Wang et al., 2024b; Ilharco et al., 2023) merge method. We refer to the model at this stage as Merge-2-Llama.
@@
Final merge Lastly, we perform a merge using the DELLA-Linear merge. With the original Llama-3.1-8B-Instruct model as the base for merging, we merge in Merge-2-Llama and Aligned-SimPO-Llama to produce the final model, Llama-SEA-LION-v3-9B-IT.
@@ 4.2 Evaluation Benchmarks
SEA Benchmarks. We evaluated the multilingual performance of each LLM using the SEA-HELM Leaderboard (Leong et al., 2023; Susanto et al., 2025)  [^5] Please visit https://leaderboard.sea-lion.ai/ for live score update of SEA-LION.. We selected SEA-HELM because the design choice of this benchmark reflects the performance of SEA culture and knowledge the most compared with other existing benchmarks DAMO-NLP-SG (2024); Lovenia et al. (2024); Wang et al. (2024a). We also evaluate on a wide-range SEA coverage language benchmark called SEACrowd Lovenia et al. (2024). This benchmark consists of all SEA languages for natural language understanding and generation datasets. However, due to maintenance reasons, we cannot reproduce the NLG benchmark of SEACrowd. Therefore, we experiment only with the NLU benchmark (zero-shot), which has 131 data subsets, 7 tasks, and 31 SEA indigenous languages.
@@ 7 Conclusion
Acknowledgment
This research is supported by the National Research Foundation, Singapore, under its National Large Language Models Funding Initiative. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.